OpenStudy (anonymous):

The polynomial functions f(x) = -13x and g(x) = -13 + x are equal.

OpenStudy (anonymous):

TRUE OR FALSE?

OpenStudy (shadowfiend):

Well the easiest way to figure this out is to check and some inputs. If x is 1, are the two equal?

OpenStudy (anonymous):

yes

OpenStudy (shadowfiend):

Are they? For x = 1, you have: f(1) = -13 * 1 g(1) = -13 + 1

OpenStudy (shadowfiend):

What is -13 * 1?

OpenStudy (anonymous):

-13

OpenStudy (shadowfiend):

And what is -13 + 1?

OpenStudy (anonymous):

-12

OpenStudy (shadowfiend):

Right. So \(f(1) \ne g(1)\), which means f(x) can't equal g(x).

OpenStudy (anonymous):

so its false.

OpenStudy (shadowfiend):

(For f(x) to be equal to g(x), the two must give the same results for all values of x.)

OpenStudy (shadowfiend):

Exactly.

OpenStudy (shadowfiend):

Also, please consider giving a medal to people who give you the right answer. It's considered polite, and helps people know they helped you :)

OpenStudy (anonymous):

O OFCOURSE I ALWAYS DO:)

OpenStudy (shadowfiend):

Thanks :):)